Assignment Score: 0/800 Resources Hint Check Answer Question 1 of 8 > How many grams of Cl are in 255 g CaCl,?
Answer #1

Molar mass of CaCl2 = 111 grams / moles

Moles of CaCl2 = mass / molar mass

= 255 / 111

= 2.2973

1 mole CaCl2 have 2 mole of Cl ( look formula )

2.2973 mole CaCl2 will have = 2*2.2973 = 4.5946 moles

Mass of Cl = mole * molar mass

= 4.5946 * 35.5

= 163.1083 grams

If you have any questions please comment

If you satisfied with the solution please rate it thanks